# Semi-definite optimization of the measured relative entropies of quantum states and channels

### v1

@@ 1.1 Background
The relative entropy [KL51] and its generalization to Rényi relative entropy [R6́1] are important distinguishability measures in information theory, finding direct operational meaning in hypothesis testing tasks [Che52, Ste, Che56, Str62, Hoe65, HK89] while being used to construct other entropic measures like mutual information and conditional entropy [Sib69, Aug78, Csi95]. There are a number of quantum generalizations of these quantities [Ume62, Pet85, Pet86, MLDS+13, WWY14], finding operational meaning in quantum hypothesis testing tasks [HP91, ON00, Nag06, ACM+07, Hay07, ANSV08, NS09] while being used to construct other entropic measures like quantum mutual information and conditional entropy (see, e.g., [KW20, Definition 7.18]). In these quantum hypothesis testing tasks, one assumes that there are many copies of the states available and furthermore that it is possible to perform a collective measurement on them. The technological capabilities required to perform such a collective measurement appear to be quite challenging, and it seems one might generally need a quantum computer to do so [DSM+22, CDCS23].
@@ 1.2 Summary of results
In this paper, we prove that the measured relative entropies of quantum states and channels can be computed by means of semi-definite optimization algorithms (also known as semi-definite programs). These algorithms have a runtime that scales efficiently with the dimension of the states and the input and output dimensions of the channels, by employing known techniques for solving semi-definite programs [PW00, AHK05, AHK12, LSW15]. Furthermore, an added benefit of these algorithms is that, not only does one obtain the optimal values of the measured relative entropies, but one also obtains numerically an optimal measurement for the measured relative entropies of states and an optimal input state and measurement for the measured relative entropies of channels. This latter capability is of significant value for applications, in which one wishes to construct a hybrid quantum-classical strategy for achieving the error rates of hypothesis testing achievable by the measured relative entropies.
@@ 3.1 Definition and basic properties
when $\alpha\in\left(0,1\right)$ or when $\alpha>1$ and $\operatorname{supp}(p)\subseteq\operatorname{supp}(q)$. Otherwise, when $\alpha>1$ and $\operatorname{supp}(p)\not\subseteq\operatorname{supp}(q)$, it is set to +$\infty$+ ∞. The Rényi relative entropy satisfies the data-processing inequality for all $\alpha\in\left(0,1\right)\cup(1,\infty)$, which means that
@@ 3.2 Variational formulas for measured Rényi relative entropy of states
$Q_{\alpha}^{M}(\rho\|\sigma)$ =$=\left\{\begin{array}[c]{cc}\inf_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]\right\}&\text{for }\alpha\in\left(0,1/2\right)\\ \inf_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega\sigma]\right\}&\text{for }\alpha\in[1/2,1)\\ \sup_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega\sigma]\right\}&\text{for }\alpha>1\end{array}\right.,$ (40)
$Q_{\alpha}^{M}(\rho\|\sigma)$ =$=\left\{\begin{array}[c]{cc}\inf_{\omega>0}\left\{\left(\operatorname{Tr}[\omega\rho]\right)^{\alpha}\left(\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]\right)^{1-\alpha}\right\}&\text{for }\alpha\in\left(0,1\right)\\ \sup_{\omega>0}\left\{\left(\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]\right)^{\alpha}\left(\operatorname{Tr}[\omega\sigma]\right)^{1-\alpha}\right\}&\text{for }\alpha>1\end{array}\right..$ (43)
@@ 4.1 Definition and basic properties
when $\operatorname{supp}(p)\subseteq\operatorname{supp}(q)$ and it is set to +$\infty$+ ∞ otherwise. It is equal to the $\alpha\rightarrow 1$ limit of the Rényi relative entropy in (19):
@@
Let us note that the convergence statement above can be made more precise by invoking (22) and [CDL+24, Lemma 23] (see also [BDSW24, Lemma 13]). Specifically, when $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$, there exists a state-dependent constant $c(\rho\|\sigma)$ such that, for all $\delta\in(0,\frac{\ln 3}{2c(\rho\|\sigma)}]$, the following bound holds:
@@ 4.3 Optimizing the measured relative entropy of states
$D^{M}(\rho\|\sigma)=\sup_{\omega,\theta>0}\left\{\operatorname{Tr}[\theta\rho]-\operatorname{Tr}[\omega\sigma]+1:\theta\leq P_{\ln}(I,\omega)\right\},$ (59)
@@
while noting that the optimal value of $\theta$ in (59) is equal to $P_{\ln}(I,\omega)$.
@@ 6.2 Optimizing the measured relative entropy of channels
$D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})=\sup_{\Omega,\Theta,\rho>0}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta\Gamma^{\mathcal{N}}]-\operatorname{Tr}[\Omega\Gamma^{\mathcal{M}}]+1:\\ \operatorname{Tr}[\rho]=1,\quad\operatorname{Tr}[H\rho]\leq E,\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\},$ (88)
@@
$=\sup_{\begin{subarray}{c}\Omega^{\prime},\Theta^{\prime}>0,\\ \rho_{R}\in\mathbb{D}(\mathcal{H}_{R}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}]-\operatorname{Tr}[\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2}]+1:\\ \Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})\end{array}\right\}$ (92)
$=\sup_{\begin{subarray}{c}\Omega^{\prime},\Theta^{\prime},\rho_{R}>0,\\ \operatorname{Tr}[\rho_{R}]=1\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\rho_{R}^{1/2}\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\rho_{R}^{1/2}\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}]+1:\\ \Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})\end{array}\right\}$ (95)
$=\sup_{\begin{subarray}{c}\Omega,\Theta,\rho_{R}>0,\\ \operatorname{Tr}[\rho_{R}]=1\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\Omega\Gamma_{RB}^{\mathcal{M}}]+1:\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\}.$ (98)
@@ 7 Conclusion
Going forward from here, we note that further work in this direction could consider combining our findings here with those of [RSB24], the latter being about measured Rényi divergences under restricted forms of measurements. We also think it is a very interesting open question, related to those from [FS17, FS23], to determine semi-definite programs for various Rényi relative entropies of quantum channels, including those based on the sandwiched [MLDS+13, WWY14] and Petz–Rényi [Pet85, Pet86] relative entropies. More generally, one could consider the same question for $\alpha$-$z$ Rényi relative entropies [AD15], and here we think the variational formulas from [MH23, Appendix B] could be useful in addressing this question.

### v2

@@ 1.1 Background
The relative entropy [KL51] and its generalization to Rényi relative entropy [R6́1] are important distinguishability measures in information theory, finding direct operational meaning in hypothesis testing tasks [Che52, Ste, Che56, Str62, Hoe65, HK89] while being used to construct other entropic measures like mutual information and conditional entropy [Sib69, Aug78, Csi95]. There are a number of quantum generalizations of these quantities [Ume62, Pet85, Pet86, MLDS+13, WWY14], finding operational meaning in quantum hypothesis testing tasks [HP91, ON00, Nag06, ACM+07, Hay07, ANSV08, NS09] while being used to construct other entropic measures like quantum mutual information and conditional entropy (see, e.g., [KW20, Definition 7.18]). In these quantum hypothesis testing tasks, one assumes that there are many copies of the states available and furthermore that it is possible to perform a collective measurement on them. The technological capabilities required to perform such a collective measurement appear to be quite challenging, and it seems one might generally need a quantum computer to do so [DSM+22, CDCS24].
@@ 1.2 Summary of results
In this paper, we prove that the measured relative entropies of quantum states and channels can be computed by means of semi-definite optimization algorithms (also known as semi-definite programs). These algorithms have runtimes that scale efficiently with the dimension of the states and the input and output dimensions of the channels, by employing known techniques for solving semi-definite programs [PW00, AHK05, AHK12, LSW15]. Furthermore, an added benefit of these algorithms is that, not only does one obtain the optimal values of the measured relative entropies, but one also obtains numerically an optimal measurement for the measured relative entropies of states and an optimal input state and measurement for the measured relative entropies of channels. This latter capability is of significant value for applications, in which one wishes to construct a hybrid quantum-classical strategy for achieving the error rates of hypothesis testing achievable by the measured relative entropies.
@@ 3.1 Definition and basic properties
when $\alpha\in\left(0,1\right)$ or when $\alpha>1$ and $\operatorname{supp}(p)\subseteq\operatorname{supp}(q)$. Otherwise, when $\alpha>1$ and $\operatorname{supp}(p)\not\subseteq\operatorname{supp}(q)$, it is set to $\infty$. The Rényi relative entropy satisfies the data-processing inequality for all $\alpha\in\left(0,1\right)\cup(1,\infty)$, which means that
@@ 3.2 Variational formulas for measured Rényi relative entropy of states
$Q_{\alpha}^{M}(\rho\|\sigma)$ $=\left\{\begin{array}[c]{cc}\inf_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]\right\}&\text{for }\alpha\in\left(0,1/2\right)\\ \inf_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega\sigma]\right\}&\text{for }\alpha\in[1/2,1)\\ \sup_{\omega>0}\left\{\alpha\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]+\left(1-\alpha\right)\operatorname{Tr}[\omega\sigma]\right\}&\text{for }\alpha>1\end{array}\right$., (40)
$Q_{\alpha}^{M}(\rho\|\sigma)$ $=\left\{\begin{array}[c]{cc}\inf_{\omega>0}\left\{\left(\operatorname{Tr}[\omega\rho]\right)^{\alpha}\left(\operatorname{Tr}[\omega^{\frac{\alpha}{\alpha-1}}\sigma]\right)^{1-\alpha}\right\}&\text{for }\alpha\in\left(0,1\right)\\ \sup_{\omega>0}\left\{\left(\operatorname{Tr}[\omega^{1-\frac{1}{\alpha}}\rho]\right)^{\alpha}\left(\operatorname{Tr}[\omega\sigma]\right)^{1-\alpha}\right\}&\text{for }\alpha>1\end{array}\right.$. (43)
@@ 4.1 Definition and basic properties
when $\operatorname{supp}(p)\subseteq\operatorname{supp}(q)$ and it is set to $\infty$ otherwise. It is equal to the $\alpha\rightarrow 1$ limit of the Rényi relative entropy in (19):
@@
Let us note that the convergence statement above can be made more precise by invoking (22) and [CDL+25, Lemma 23] (see also [BDSW24, Lemma 13]). Specifically, when $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$, there exists a state-dependent constant $c(\rho\|\sigma)$ such that, for all $\delta\in(0,\frac{\ln 3}{2c(\rho\|\sigma)}]$, the following bound holds:
@@ 4.3 Optimizing the measured relative entropy of states
$D^{M}(\rho\|\sigma)=\sup_{\omega>0,\theta\in\operatorname{Herm}}\left\{\operatorname{Tr}[\theta\rho]-\operatorname{Tr}[\omega\sigma]+1:\theta\leq P_{\ln}(I,\omega)\right\},$ (59)
@@
while noting that the optimal value of $\theta$ in (59) is equal to $P_{\ln}(I,\omega)$, the latter being a Hermitian operator.
@@ 6.2 Optimizing the measured relative entropy of channels
$D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})=\sup_{\Omega,\rho>0,\Theta\in\operatorname{Herm}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta\Gamma^{\mathcal{N}}]-\operatorname{Tr}[\Omega\Gamma^{\mathcal{M}}]+1:\\ \operatorname{Tr}[\rho]=1,\quad\operatorname{Tr}[H\rho]\leq E,\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\},$ (88)
@@
$=\sup_{\begin{subarray}{c}\Omega^{\prime}>0,\Theta^{\prime}\in\operatorname{Herm},\\ \rho_{R}\in\mathbb{D}(\mathcal{H}_{R}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}]-\operatorname{Tr}[\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2}]+1:\\ \Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})\end{array}\right\}$ (92)
$=\sup_{\begin{subarray}{c}\Omega^{\prime},\rho_{R}>0,\Theta^{\prime}\in\operatorname{Herm}\\ \operatorname{Tr}[\rho_{R}]=1\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\rho_{R}^{1/2}\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\rho_{R}^{1/2}\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}]+1:\\ \Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})\end{array}\right\}$ (95)
$=\sup_{\begin{subarray}{c}\Omega,\rho_{R}>0,\Theta\in\operatorname{Herm}\\ \operatorname{Tr}[\rho_{R}]=1\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\Omega\Gamma_{RB}^{\mathcal{M}}]+1:\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\}.$ (98)
@@
$\Omega^{\prime}>0,\quad\Theta^{\prime}\in\operatorname{Herm}\quad\Leftrightarrow\quad\Omega>0,\quad\Theta\in\operatorname{Herm},$ (101)
@@ 7 Conclusion
Going forward from here, we note that further work in this direction could consider combining our findings here with those of [RSB24], the latter being about measured Rényi divergences under restricted forms of measurements. We also think it is a very interesting open question, related to those from [FS17, FS23], to determine semi-definite programs for various Rényi relative entropies of quantum channels, including those based on the sandwiched [MLDS+13, WWY14] and Petz–Rényi [Pet85, Pet86]  relative entropies. More generally, one could consider the same question for $\alpha$-$z$ Rényi relative entropies [AD15], and here we think the variational formulas from [MH24, Appendix B] could be useful in addressing this question.